When a monopolist lowers a price from $80 to $70, the quantity that the firm is able to sell increases from 100 to 150. The change in revenue associated with the price effect is equal to
a. $3,500. d. −$1,000.
b. −$3,500. e. $4,000.

c. $1,000.

please explain
Answer #1

change in revenue due to price effect = (70 - 80)* 100 = -10 * 100 = -1000

option d is the correct asnwer